OpenStudy (anonymous):

Charlie runs at the speed of 3 yards per second. About how many miles per hour does charlie run??

OpenStudy (mathstudent55):

How many seconds are there in an hour?

OpenStudy (anonymous):

360?

OpenStudy (anonymous):

6.136 mph

OpenStudy (anonymous):

thank you

OpenStudy (mathstudent55):

\(1 ~hour \times \dfrac{60~min}{1~hour} \times \dfrac{60~sec}{1~min}\) \(=1 ~\cancel{hour} \times \dfrac{60~\cancel{min}}{1~\cancel{hour}} \times \dfrac{60~sec}{1~\cancel{min}} = 3600 ~sec\) 1 hour = 3600 sec

OpenStudy (mathstudent55):

He runs 3 yards in 1 second. There are 3600 seconds in 1 hour, so he runs 3 * 3600 yards in 1 hour 3 * 3600 = 10,800 His speed is 10,800 yd/hr There are 1760 yards in a mile, so his speed is: 10,800/1760 mph = 6.14 mph

OpenStudy (mathstudent55):

This entire conversion can be done in one calculation: \(3\dfrac{yd}{sec} \times \dfrac{1~mile}{1760~yd} \times \dfrac{60~sec}{1~min} \times \dfrac{60~min}{1~hour} \) \(= 6.14 ~mph\)